# April 3, 2025

### v1

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
Since the right-hand sides of the estimates in (7), (9) and (11) are independent of $m$ and $\kappa$, we obtain that the following convergences hold up to passing to a subsequence:
@@
[…]
$\dot{\bm{u}}_{\kappa}$ $\rightharpoonup\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{V}(\Omega))\textup{ as }\kappa\to 0,$
[…]
The convergence $\bm{u}_{\kappa}\rightharpoonup\bm{u}$ in $\mathcal{C}^{0}([0,T];\bm{L}^{2}(\Omega))$ as $\kappa\to 0$ holds thanks to Theorem 10.1.20 in [34] (see also [18]). The weak convergences of the sequences $\{\dot{\bm{u}}_{\kappa}\}_{\kappa>0}$ and $\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}$ and the assumed admissibility of the applied body force $\bm{f}$ in turn imply that:
$\dot{\bm{u}}_{\kappa}(T)\rightharpoonup\dot{\bm{u}}(T),\quad\textup{ in }\bm{V}^{\ast}(\Omega)\textup{ as }\kappa\to 0.$
Let us now observe that an application of (9) gives
$\int_{0}^{T}\chi_{(0,T_{0})}\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t$
$=\int_{0}^{T}\int_{\Gamma}\chi_{(0,T_{0})}\left|\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\}^{-}\right|^{2}\,\mathrm{d}\Gamma\,\mathrm{d}t+\int_{0}^{T}\int_{\Gamma}\chi_{(0,T_{0})}\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\}^{-}(\bm{I}\cdot\bm{q})\,\mathrm{d}\Gamma\,\mathrm{d}t\to 0,$
as $\kappa\to 0$. The Lipschitz continuity and monotonicity of the operator $\tilde{\mathcal{N}}$ introduced in (1) put us in position to apply Theorem 9.13-2 in [10] so as to infer that
$\{[\bm{I}+u_{i}\bm{e}^{i}]\cdot\bm{q}\}^{-}=0,\quad\textup{ in }L^{2}(T_{0},T;\bm{V}^{\ast}(\Omega)),$
as well as the validity of the fifth convergence in (17). In particular, the latter implies that $\bm{u}\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$. Clearly, the assumed admissibility of the applied body forces (Definition 3.1) gives that the constraint is inactive in $(T_{0},T)$. This completes the proof. ∎
Let us also recall a result on vector-valued measures (cf., e.g., page 380 of [13]).
Theorem 4.3 (Dinculeanu-Zinger).
Let $\omega$ be a compact Hausdorff space and let $X$ be a Banach space satisfying the Radon-Nikodym property. Let $\mathcal{F}$ be the collection of Borel sets of $\omega$.
There exists an isomorphism between $(\mathcal{C}^{0}(\omega;X))^{\ast}$ and the space of the regular Borel measures with finite variation taking values in $X^{\ast}$. In particular, for each $F\in(\mathcal{C}^{0}(\omega;X))^{\ast}$, there exists a unique regular Borel measure $\mu:\mathcal{F}\to X^{\ast}$ in $\mathcal{M}(\omega;X^{\ast})$ with finite variation such that
$\langle\langle\alpha,F\rangle\rangle_{X^{\ast},X}=\int_{\omega}\langle\,\mathrm{d}\mu(t),\alpha(t)\rangle_{X^{\ast},X}\,\mathrm{d}t,\quad\textup{ for all }\alpha\in\mathcal{C}^{0}(\omega;X).$
The next lemma aims to verify that the limit $\bm{u}$ introduced in Lemma 4.4 satisfies the initial condition for the velocity. This property hinges on the fact that the confinement condition we are here employing is of Signorini type.
Lemma 4.5.
The limit $\bm{u}$ recovered in (17) satisfies the following properties:
  • $(i)$
    $\dot{\bm{u}}(0)=\bm{u}_{1}$;
  • $(ii)$
    $\ddot{\bm{u}}\in\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega))$.

### v3

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
Testing the variational equations of Problem $\mathcal{P}_{\kappa}(\Omega)$ at $\dot{\bm{u}}_{\kappa}(t)$ for a.a. $t\in(0,T)$, and proceeding similarly to item (ii) in the proof of Theorem 4.3 we obtain energy estimates similar to (13), so that:
(24) $\{\bm{u}_{\kappa}\}_{\kappa>0}$ $\textup{ is bounded in }L^{\infty}(0,T;\bm{V}(\Omega))\textup{ independently of }\kappa,$
$\{\dot{\bm{u}}_{\kappa}\}_{\kappa>0}$ $\textup{ is bounded in }L^{\infty}(0,T;\bm{L}^{2}(\Omega))\cap L^{2}(0,T;\bm{V}(\Omega))\textup{ independently of }\kappa,$
$\left\|\left\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(0,T;L^{2}(\Gamma))}\leq\sqrt{\kappa Tbe^{T}}.$
Therefore, the following convergences hold up to passing to a subsequence indexed over $\kappa_{n}$, where $\{\kappa_{n}\}_{n=1}^{\infty}$ denotes a sequence of positive real numbers such that $\kappa_{n}\to 0^{+}$ as $n\to\infty$:
$\bm{u}_{\kappa_{n}}$ $\overset{\ast}{\rightharpoonup}\bm{u},\quad\textup{ in }L^{\infty}(0,T;\bm{V}(\Omega))\textup{ as }n\to\infty,$
$\dot{\bm{u}}_{\kappa_{n}}$ $\overset{\ast}{\rightharpoonup}\dot{\bm{u}},\quad\textup{ in }L^{\infty}(0,T;\bm{L}^{2}(\Omega))\textup{ as }n\to\infty,$
$\dot{\bm{u}}_{\kappa_{n}}$ $\rightharpoonup\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{V}(\Omega))\textup{ as }n\to\infty.$
The convergence $\bm{u}_{\kappa_{n}}\rightharpoonup\bm{u}$ in $\mathcal{C}^{0}([0,T];\bm{V}(\Omega))$ as $n\to\infty$ holds thanks to Theorem 10.1.20 in [43] (see also [25]). The convergence $\bm{u}_{\kappa_{n}}\to\bm{u}$ in $\mathcal{C}^{0}([0,T];\bm{L}^{2}(\Omega))$ as $n\to\infty$ holds thanks to the Aubin-Lions-Simon theorem (Theorem 4.2). Let us now observe that an application of (15) gives
$-\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}u_{i,\kappa_{n}}(t)\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma\,\mathrm{d}t$
$=-\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\left([\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\right)\,\mathrm{d}\Gamma\,\mathrm{d}t+\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}(\bm{I}\cdot\bm{q})\,\mathrm{d}\Gamma\,\mathrm{d}t$
$=-\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{+}\,\mathrm{d}\Gamma\,\mathrm{d}t$
$\quad+\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\,\mathrm{d}\Gamma\,\mathrm{d}t+\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}(\bm{I}\cdot\bm{q})\,\mathrm{d}\Gamma\,\mathrm{d}t$
$=\int_{0}^{T}\int_{\Gamma}\left|\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\right|^{2}\,\mathrm{d}\Gamma\,\mathrm{d}t+\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa_{n}}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}(\bm{I}\cdot\bm{q})\,\mathrm{d}\Gamma\,\mathrm{d}t\to 0,$
as $n\to\infty$. The Lipschitz continuity and monotonicity of the operator $\{\cdot\}^{-}$ (Lemma 4.1) put us in position to apply the corollary to the Minty-Browder theorem (Theorem 12.5-2 in [15]) so as to infer that
$\{[\bm{I}+u_{i}\bm{e}^{i}]\cdot\bm{q}\}^{-}=0,\quad\textup{ in }L^{2}(0,T;L^{2}(\Gamma)),$
as well as the validity of the fourth convergence in (23). In particular, the convergence $\bm{u}_{\kappa_{n}}\rightharpoonup\bm{u}$ in $\mathcal{C}^{0}([0,T];\bm{V}(\Omega))$ as $n\to\infty$ obtained beforehand as a result of Theorem 10.1.20 of [43] implies that $\bm{u}\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$ and that $\bm{u}(0)=\bm{u}_{0}$. This completes the proof. ∎
The next lemma aims to verify that the limit $\bm{u}$ introduced in Lemma 4.5 satisfies the initial condition for the velocity. Due to the limited regularity of the limit $\ddot{\bm{u}}$, we need to restrict our analysis to applied body forces $\bm{f}=(f^{i})\in H^{1}(0,T;\bm{L}^{2}(\Omega))$.
Lemma 4.6.
Assume that $\bm{f}=(f^{i})\in H^{1}(0,T;\bm{L}^{2}(\Omega))$. Then the limit $\dot{\bm{u}}\in\mathcal{C}^{0}([0,T];\bm{L}^{2}(\Omega))$ and is such that $\dot{\bm{u}}(0)=\bm{u}_{1}$ and $\dot{\bm{u}}(T)\in\bm{L}^{2}(\Omega)$. Besides, up to passing to a suitable subsequence $\{\dot{\bm{u}}_{\kappa_{n}}\}_{n=1}^{\infty}$, where $\{\kappa_{n}\}_{n=1}^{\infty}$ is a sequence such that $\kappa_{n}\to 0^{+}$ as $n\to\infty$, it results that:
$\dot{\bm{u}}_{\kappa_{n}}\to\dot{\bm{u}}$ , $,\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Omega))\textup{ as }n\to\infty,$
$\dot{\bm{u}}_{\kappa_{n}}(T)\rightharpoonup\dot{\bm{u}}(T)$ , $,\quad\textup{ in }\bm{L}^{2}(\Omega)\textup{ as }n\to\infty.$
Proof.
For $a.$ $t\in(0,T)$, and for all $\kappa>0$, any solution $\bm{u}_{\kappa}$ for Problem $\mathcal{P}_{\kappa}(\Omega)$ satisfies the following equation:
(25) $2\rho\ddot{\bm{u}}_{\kappa}(t)+(\tilde{\mathcal{A}}\bm{u}_{\kappa})(t)+(\tilde{\mathcal{B}}\dot{\bm{u}}_{\kappa})(t)+\dfrac{1}{\kappa}(\tilde{\mathcal{N}}\bm{u}_{\kappa})(t)=\bm{f}(t),\quad\textup{ in }\bm{V}^{\ast}(\Omega).$
Testing (25) along any element $\bm{v}\in\bm{H}^{1}_{0}(\Omega)$, we obtain that the penalty term vanishes. Therefore, the function $\bm{u}_{\kappa}$ satisfies the following equation
in the sense of distributions in $(0,T)$. Note that (26) is a linear equation; this implies that
$\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}\textup{ is bounded in }L^{2}(0,T;\bm{H}^{-1}(\Omega))\textup{ independently of }\kappa,$
so that, up to passing to a suitable subsequence $\{\ddot{\bm{u}}_{\kappa_{n}}\}_{n=1}^{\infty}$ where $\{\kappa_{n}\}_{n=1}^{\infty}$ is a sequence such that $\kappa_{n}\to 0^{+}$ as $n\to\infty$ we obtain:
$\ddot{\bm{u}}_{\kappa_{n}}\rightharpoonup\ddot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{H}^{-1}(\Omega))\textup{ as }n\to\infty.$
Therefore, it is immediate to observe that $\dot{\bm{u}}_{\kappa_{n}}\rightharpoonup\dot{\bm{u}}$ in $\mathcal{C}^{0}([0,T];\bm{H}^{-1}(\Omega))$ as $n\to\infty$ and that:
(27) $\dot{\bm{u}}_{\kappa_{n}}(T)\rightharpoonup\dot{\bm{u}}(T),\quad\textup{ in }\bm{H}^{-1}(\Omega)\textup{ as }n\to\infty.$
The latter allows us to show that $\dot{\bm{u}}(0)=\bm{u}_{1}$ and $\dot{\bm{u}}(T)\in\bm{H}^{-1}(\Omega)$. Additionally, an application of the the Aubin-Lions-Simon theorem (Theorem 4.2) gives:
$\dot{\bm{u}}_{\kappa_{n}}\to\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Omega))\textup{ as }n\to\infty.$
For each $\kappa>0$ and for a.a. $0\leq t\leq T$, define the energy:
(28) $E_{\kappa}(t)=\rho\int_{\Omega}|\dot{\bm{u}}_{\kappa}(t)|^{2}\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{u}_{\kappa}(t))\,\mathrm{d}x$
$\quad+\dfrac{1}{2\kappa}\int_{\Gamma}|\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}|^{2}\,\mathrm{d}\Gamma-\int_{\Omega}f^{i}(t)u_{i,\kappa}(t)\,\mathrm{d}x.$
Observe that:
(29) $\int_{0}^{T}|E_{\kappa}(t)|\,\mathrm{d}t\leq\rho\|\dot{\bm{u}}_{\kappa}(t)\|_{\bm{L}^{2}(\Omega)}^{2}+\dfrac{\max_{i,j,k,\ell}\|A^{ijk\ell}\|_{\mathcal{C}^{0}(\overline{\Omega})}c_{0}^{2}}{2}\int_{0}^{T}\|\bm{u}_{\kappa}(t)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}t$
$\quad+\dfrac{1}{2\kappa}\int_{\Gamma}\left|\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\right|^{2}\,\mathrm{d}\Gamma+\dfrac{1}{2}\int_{0}^{T}\left\{\sum_{i=1}^{3}\|f^{i}(t)\|_{L^{2}(\Omega)}^{2}\right\}\,\mathrm{d}t$
$\quad+\dfrac{1}{2}\int_{0}^{T}\left\{\sum_{i=1}^{3}\|u_{i,\kappa}(t)\|_{L^{2}(\Omega)}^{2}\right\}\,\mathrm{d}t.$
Since the right-hand side of (29) is bounded independently of $\kappa$ thanks to (24), we infer that $E_{\kappa}\in L^{1}(0,T)$, and that the family $\{E_{\kappa}\}_{\kappa>0}$ is bounded in $L^{1}(0,T)$ independently of $\kappa$. Observe that, thanks to Corollary 10.1.26 of [43], it results that $E_{\kappa}$ is differentiable a.e. in $(0,T)$ and that $\dot{E}_{\kappa}\in L^{1}(0,T)$. Therefore, we obtain that $E_{\kappa}\in W^{1,1}(0,T)\subset{{}}{}$.
Differentiating $E_{\kappa}$ and invoking the variational equations of Problem $\mathcal{P}_{\kappa}(\Omega)$ tested at $\bm{v}=\bm{u}_{\kappa}(t)$ give:
(30) $\dot{E}_{\kappa}(t)=\rho\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Omega}|\dot{\bm{u}}_{\kappa}(t)|^{2}\,\mathrm{d}x\right)+\dfrac{1}{2}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{u}_{\kappa}(t))\,\mathrm{d}x\right)$
$\quad+\dfrac{1}{2\kappa}\int_{\Gamma}\left|\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\right|^{2}\,\mathrm{d}\Gamma-\int_{\Omega}f^{i}(t)\dot{u}_{i,\kappa}(t)\,\mathrm{d}x-\int_{\Omega}\dot{f}^{i}(t)u_{i,\kappa}(t)\,\mathrm{d}x$
$=-\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}_{\kappa}(t))\,\mathrm{d}x-\int_{\Omega}\dot{f}^{i}(t)u_{i,\kappa}(t)\,\mathrm{d}x.$
Taking the absolute value in (30) and employing Young’s inequality [57] gives:
(31) $|\dot{E}_{\kappa}(t)|\leq\left|\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}_{\kappa}(t))\,\mathrm{d}x\right|+\dfrac{1}{2}\sum_{i=1}^{3}\|\dot{f}^{i}(t)\|_{L^{2}(\Omega)}^{2}+\dfrac{1}{2}\sum_{i=1}^{3}\|u_{i,\kappa}(t)\|_{L^{2}(\Omega)}^{2}.$
Observe that, thanks to (24), the right-hand side of (31) is bounded independently of $\kappa$ showing that the family $\{\dot{E}_{\kappa}\}_{\kappa>0}$ is bounded in $L^{1}(0,T)$ independently of $\kappa$.
The continuity of $E_{\kappa}$ in $[0,T]$ and the continuity of the second, third and fourth term on the right-hand side of (28) implies that the mapping
(32) $0,T]\ni t\mapsto\int_{\Omega}|\dot{\bm{u}}_{\kappa}(t)|^{2}\,\mathrm{d}x,$
is continuous and, thanks to (23), is bounded in $[0,T]$ independently of $\kappa$. The boundedness of the mapping in (32) in $[0,T]$ independently of $\kappa$ is equivalent to stating that there exists a constant $C>0$ independent of $\kappa$ such that
$\max_{t\in[0,T]}\|\dot{\bm{u}}_{\kappa}(t)\|_{\bm{L}^{2}(\Omega)}\leq C,\quad\textup{ for all }\kappa>0,$
and, in particular, that $\{\dot{\bm{u}}_{\kappa}(T)\}_{\kappa>0}$ is bounded in $\bm{L}^{2}(\Omega)$ independently of $\kappa$. Therefore, up to passing to a suitable subsequence $\{\dot{\bm{u}}_{\kappa_{n}}(T)\}_{n=1}^{\infty}$ where $\{\kappa_{n}\}_{n=1}^{\infty}$ is a sequence such that $\kappa_{n}\to 0^{+}$ as $n\to\infty$ we obtain:
$\dot{\bm{u}}_{\kappa_{n}}(T)\rightharpoonup\bm{U}_{T},\quad\textup{ in }\bm{L}^{2}(\Omega)\textup{ as }n\to\infty.$
Since we have already observed in (27) that $\dot{\bm{u}}_{\kappa_{n}}(T)\rightharpoonup\dot{\bm{u}}(T)$ in $\bm{H}^{-1}(\Omega)$ as $n\to\infty$, the uniqueness of the weak limit gives that $\dot{\bm{u}}(T)=\bm{U}_{T}\in\bm{L}^{2}(\Omega)$. ∎
We are now ready to establish the main result of this paper: the existence of solutions for Problem $\mathcal{P}(\Omega)$.
Theorem $4.$
Assume that $\bm{f}\in H^{1}(0,T;\bm{L}^{2}(\Omega))$. Then Problem $\mathcal{P}(\Omega)$ admits at least one solution.
Proof.
Fix $\bm{v}=(v_{i})\in L^{\infty}(0,T;\bm{U}(\Omega))$ such that $\dot{\bm{v}}\in L^{\infty}(0,T;\bm{L}^{2}(\Omega))\cap L^{2}(0,T;\bm{V}(\Omega))$. For any given $t\in[0,T]$, let us test equation (25) at the element $(\bm{v}(t)-\bm{u}_{\kappa}(t))\in\bm{V}(\Omega)$. Thanks to the convergence $\dot{\bm{u}}_{\kappa_{n}}(T)\rightharpoonup\dot{\bm{u}}(T)$ in $\bm{L}^{2}(\Omega)$ as $n\to\infty$ established in Lemma 4.6, and thanks to the convergence $\dot{\bm{u}}_{\kappa_{n}}\rightharpoonup\dot{\bm{u}}$ in $L^{2}(0,T;\bm{L}^{2}(\Omega))$ as $n\to\infty$ established in Lemma 4.5, we obtain that:
(33) $\int_{0}^{T}\langle\ddot{\bm{u}}_{\kappa_{n}}(t),\bm{v}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t$
$=\int_{\Omega}\dot{\bm{u}}_{\kappa_{n}}(T)\cdot\bm{v}(T)\,\mathrm{d}x-\int_{\Omega}\bm{u}_{1}\cdot\bm{v}(0)\,\mathrm{d}x-\int_{0}^{T}\int_{\Omega}\dot{\bm{u}}_{\kappa_{n}}(t)\cdot\dot{\bm{v}}(t)\,\mathrm{d}x\,\mathrm{d}t$
$\to\int_{\Omega}\dot{\bm{u}}(T)\cdot\bm{v}(T)\,\mathrm{d}x-\int_{\Omega}\bm{u}_{1}\cdot\bm{v}(0)\,\mathrm{d}x-\int_{0}^{T}\int_{\Omega}\dot{\bm{u}}(t)\cdot\dot{\bm{v}}(t)\,\mathrm{d}x\,\mathrm{d}t,\quad\textup{ as }n\to\infty.$
The convergence $\dot{\bm{u}}_{\kappa_{n}}(T)\rightharpoonup\dot{\bm{u}}(T)$ in $\bm{L}^{2}(\Omega)$ as $n\to\infty$ established in Lemma 4.6, the convergence $\dot{\bm{u}}_{\kappa_{n}}\to\dot{\bm{u}}$ in $L^{2}(0,T;\bm{L}^{2}(\Omega))$ as $n\to\infty$ established in Lemma 4.6, and the convergence $\bm{u}_{\kappa_{n}}\to\bm{u}$ in $\mathcal{C}^{0}([0,T];\bm{L}^{2}(\Omega))$ as $n\to\infty$ established in Lemma 4.5 give:
(34) $-\int_{0}^{T}\langle\ddot{\bm{u}}_{\kappa_{n}}(t),\bm{u}_{\kappa_{n}}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t$
$=-\int_{\Omega}\dot{\bm{u}}_{\kappa_{n}}(T)\cdot\bm{u}_{\kappa_{n}}(T)\,\mathrm{d}x+\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+\|\dot{\bm{u}}_{\kappa_{n}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}$
$\to-\int_{\Omega}\dot{\bm{u}}(T)\cdot\bm{u}(T)\,\mathrm{d}x+\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+\|\dot{\bm{u}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2},\quad\textup{ as }n\to\infty.$
The convergence $\dot{\bm{u}}_{\kappa_{n}}\rightharpoonup\dot{\bm{u}}$ in $L^{2}(0,T;\bm{V}(\Omega))$ as $n\to\infty$ established in (23) and the continuity of the operator $\mathcal{B}$ (Lemma 4.4) imply that:
(35) $\int_{0}^{T}\langle\mathcal{B}\dot{\bm{u}}_{\kappa_{n}}(t),\bm{v}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\to\int_{0}^{T}\langle\mathcal{B}\dot{\bm{u}}(t),\bm{v}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t,\quad\textup{ as }n\to\infty.$
The convergence $\bm{u}_{\kappa_{n}}\rightharpoonup\bm{u}$ in $\mathcal{C}^{0}([0,T];\bm{V}(\Omega))$ as $n\to\infty$ established in (23) implies that:
(36) $\limsup_{n\to\infty}\left(-\int_{0}^{T}\langle\mathcal{B}\dot{\bm{u}}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)$
$\leq-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}(T))e_{i\|j}(\bm{u}(T))\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x.$
The convergence $\bm{u}_{\kappa_{n}}\rightharpoonup\bm{u}$ in $L^{2}(0,T;\bm{V}(\Omega))$ as $n\to\infty$ established in (23) and the continuity of the operator $\mathcal{A}$ (Lemma 4.3) imply that:
(37) $\int_{0}^{T}\langle\mathcal{A}\bm{u}_{\kappa_{n}}(t),\bm{v}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\to\int_{0}^{T}\langle\mathcal{A}\bm{u}(t),\bm{v}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t,\quad\textup{ as }n\to\infty.$
The convergence $\bm{u}_{\kappa_{n}}\rightharpoonup\bm{u}$ in $L^{2}(0,T;\bm{V}(\Omega))$ as $n\to\infty$ established in (23) implies that:
(38) $\limsup_{n\to\infty}\left(-\int_{0}^{T}\langle\mathcal{A}\bm{u}_{\kappa_{n}}(t),\bm{u}_{\kappa_{n}}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)\leq-\int_{0}^{T}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}(t))e_{i\|j}(\bm{u}(t))\,\mathrm{d}x\,\mathrm{d}t.$
For what concerns the penalty term, we observe that the monotonicity of $\{\cdot\}^{-}$ (Lemma 4.1) gives
(39) $\dfrac{1}{\kappa}\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\left([\bm{I}+v_{i}(t)\bm{e}^{i}]\cdot\bm{q}-[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right)\,\mathrm{d}\Gamma\,\mathrm{d}t\leq 0,$
for all $\kappa>0$.
Finally, the convergence $\bm{u}_{\kappa_{n}}\rightharpoonup\bm{u}$ in $L^{2}(0,T;\bm{L}^{2}(\Omega))$ as $n\to\infty$ established in (23) implies that:
@@
$2\rho\int_{\Omega}\dot{\bm{u}}(T)\cdot(\bm{v}(T)-\bm{u}(T))\,\mathrm{d}x-2\rho\int_{\Omega}\bm{u}_{1}\cdot(\bm{v}(0)-\bm{u}_{0})\,\mathrm{d}x-2\rho\int_{0}^{T}\int_{\Omega}\dot{\bm{u}}(t)\cdot(\dot{\bm{v}}(t)-\dot{\bm{u}}(t))\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}(T))e_{i\|j}(\bm{u}(T))\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x$
which is exactly the governing model of Problem $\mathcal{P}(\Omega)$. This completes the proof. ∎
Finally, we observe that if we assume that there exists a shut-down time $T_{0}>0$ such that $\bm{f}(t)=\bm{0}$ for a.a. $t>T_{0}$, we are able to establish that, for certain materials, the linearly elastic viscoelastic body returns to its reference configuration when the applied body force is lifted and that, in agreement with the theory of Kelvin-Voigt materials [11, 52], the decay is exponential.
Theorem 4.5.
Assume that there exists $0<T_{0}<T$ such that $\bm{f}(t)=\bm{0}$ for a.a. $T_{0}<t<T$. Assume that the following smallness condition holds:
(41) $\dfrac{\rho}{C_{e}^{(1)}\max_{i,j,k,\ell}\|B^{ijk\ell}\|_{\mathcal{C}^{0}(\overline{\Omega})}}\leq\dfrac{1}{2c_{0}^{2}C_{e}^{(2)}}.$
For each $\kappa>0$ and for a.a. $T_{0}\leq t\leq T$, define the energy:
$E_{\kappa}(t):=\rho\int_{\Omega}|\dot{\bm{u}}_{\kappa}(t)|^{2}\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{u}_{\kappa}(t))\,\mathrm{d}x+\dfrac{1}{2\kappa}\int_{\Gamma}|\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}|^{2}\,\mathrm{d}\Gamma.$
Then, $E_{\kappa}\in AC([T_{0},T])$ and there exists a constant $\tilde{C}=\tilde{C}(\Omega,\lambda,\mu,\xi,\theta,\rho)>0$ and a constant $c_{{\textup{d}}}=c_{{\textup{d}}}(\Omega,\lambda,\mu,\xi,\theta,\rho)>0$ such that:
(42) $\|\bm{u}_{\kappa}(t)\|_{\bm{V}(\Omega)}+\|\dot{\bm{u}}_{\kappa}(t)\|_{\bm{L}^{2}(\Omega)}\leq\sqrt{\tilde{C}}\exp\left(-\dfrac{c_{{\textup{d}}}}{2}(t-T_{0})\right),$
for a.a. $T_{0}<t<T$ and all $\kappa>0$.